# Thunder-NUBench: A Benchmark for LLMs’ Sentence-Level Negation Understanding

### v1

@@ Appendix I Models We Use in Experiments
Listing 9: ScoNe/ScoNe.yaml
@@
3
4def process_docs(dataset: datasets.Dataset) -> datasets.Dataset:
@@
7
8 choices = ["Yes", "No"]
9 if doc[’gold_label_edited’] == "entailment":
10 gold = 0
@@
Listing 10: ScoNe/utils.py
Appendix L Finetuning via LLaMA-Factory
We detail our supervised fine-tuning setup using LLaMA-Factory (Zheng et al., 2024) with LoRA (Hu et al., 2022) on Thunder-NUBench training data, including configuration of the fine-tuning and instruction-based examples in Alpaca format (Taori et al., 2023).
The YAML configuration provided in Listing 11 is specific to the LLaMA-3.1-8B model. Other models (e.g., Qwen or Mistral) can be fine-tuned similarly by modifying the model_name_or_path and template fields in the configuration file accordingly.
@@
Listing 11: Llama-3.1-8B_lora_sft.yaml
@@
3 "instruction": "Logically negate the sentence below. If the sentence includes A and B, use not A or not B’. If it includes A or B, use not A and not B’. Also apply not or use complementary antonyms on the main verb(s) of the entire sentence.",
[…]
@@
8 "instruction": "Logically negate the sentence below. If the sentence includes A and B, use not A or not B’. If it includes A or B, use not A and not B’. Also apply not or use complementary antonyms on the main verb(s) of the entire sentence.",
[…]
@@
Listing 12: Sentence-Negation Pair dataset for training in alpaca format.
We use 4 NVIDIA RTX 3090 GPUs (24GB) with CUDA 12.4 for all model training and evaluation. Both supervised fine-tuning and evaluation are completed in under an hour using these resources.
default prompt instruction prompt reasoning prompt average
zeroshot
fewshot
@@
3B-4B Llama Llama-3.2-3B acc 0.475
0.603
@@
0.616
@@
0.610 0.734
average acc 0.455 0.639 0.496 0.647 0.541 0.656 0.497 0.647
acc_
norm
0.477 0.653 0.516 0.660 0.560 0.671 0.518 0.662
default prompt instruction prompt reasoning prompt average
zeroshot
fewshot
(±SD)
zeroshot
fewshot
(±SD)
zeroshot
fewshot
(±SD)
zeroshot
fewshot
(±SD)
7B-8B Llama Llama-3.1-8B acc 0.464
0.662
(±0.007)
0.477
0.643
@@
0.464
0.656
(±0.007)
0.468 0.654
acc_
norm
0.487
0.676
(±0.010)
0.504
0.657
(±0.008)
0.493
0.673
(±0.007)
0.495 0.669
Llama3.1-8B- Instruct acc 0.492
0.706
(±0.004)
0.564
0.698
(±0.007)
0.513
0.692
(±0.005)
0.523 0.699
acc_
norm
0.518
0.718
(±0.005)
0.584
0.706
(±0.006)
0.535
0.703
(±0.006)
0.546 0.709
gemma gemma-7b acc 0.471
0.648
(±0.006)
0.492
0.621
(±0.005)
0.504
0.646
(±0.002)
0.489 0.638
acc_
norm
0.495
0.660
(±0.006)
0.518
0.635
(±0.006)
0.530
0.658
(±0.003)
0.514 0.651
gemma-7b- it acc 0.559
0.731

### v3

@@ 8 Paraphrased sentence:
13 aggregation: mean
14 higher_is_better: true
Listing 7: NUBench/NUBench_completion.yaml
1task: nubench_option
2dataset_path: {dataset_path}/NUBench
3dataset_name: multiple-choice
4output_type: generate_until
5test_split: test
6fewshot_split: demonstration
7process_docs: !function utils.process_docs_option
8doc_to_text: "{{query}}"
9doc_to_target: "{{answerLetter}}"
10generation_kwargs:
11 until:
12 - "</s>"
[…]
14metric_list:
15 - metric: exact_match
16 aggregation: mean
17 higher_is_better: true
18 ignore_punctuation: true
19 ignore_case: true
20filter_list:
21 - name: get_response
22 filter:
23 - function: "regex"
[…]
25 - function: remove_whitespace
26 - function: "regex"
[…]
28 - function: take_first
29dataset_kwargs:
30 trust_remote_code: true
Listing 8: NUBench/NUBench_option.yaml
@@
3import random
4
5def process_docs_completion(dataset: datasets.Dataset) -> datasets.Dataset:
@@
8
9 if doc.get("choice2_type", "") == "non-applicable":
10 choices = [doc["choice1"], doc["choice3"], doc["choice4"]]
@@
20
21
22def process_docs_option(dataset: datasets.Dataset, seed: int = 42) -> datasets.Dataset:
23 rng = random.Random(seed)
24
25 def _process_doc(doc):
[…]
27
28 c1 = doc.get("choice1", "")
29 c2 = doc.get("choice2", "")
30 c3 = doc.get("choice3", "")
31 c4 = doc.get("choice4", "")
32 c2_na = str(doc.get("choice2_type", "")).lower() == "non-applicable"
33
34 gathered = []
35 keys = []
36
37 if c1:
38 gathered.append(c1); keys.append("choice1")
39 if c2 and not c2_na:
40 gathered.append(c2); keys.append("choice2")
41 if c3:
42 gathered.append(c3); keys.append("choice3")
43 if c4:
44 gathered.append(c4); keys.append("choice4")
45
46 assert len(gathered) >= 3, "Need at least 3 choices for multiple-choice."
47
48 paired = list(zip(gathered, keys))
49 rng.shuffle(paired)
50 choices, choice_keys = zip(*paired)
51
52 try:
53 answer_idx = choice_keys.index("choice1")
54 except ValueError:
55 answer_idx = 0
56
57 label_map = ["A", "B", "C", "D"]
58 labels = label_map[:len(choices)]
59
60 lines = [
61 "Given the following instruction and candidate answers, choose the single best answer.",
62 f"Instruction: {initial_prompt}",
63 ]
64 for lab, ch in zip(labels, choices):
65 lines.append(f"{lab}. {ch}")
66 joined = ", ".join(labels)
67 lines += ["", f"Your response should be one of {joined}.", "Only output the letter.", "Answer:"]
[…]
69
70 return {
71 "query": prompt,
72 "choices": list(choices),
73 "choice_keys": list(choice_keys),
74 "answerKey": answer_idx,
75 "answerLetter": labels[answer_idx],
76 }
77
78 return dataset.map(_process_doc)
Listing 9: NUBench/utils.py
Appendix N Finetuning via LLaMA-Factory
We detail our supervised fine-tuning setup using LLaMA-Factory (zheng2024llamafactory) with LoRA (hu2022lora) on Thunder-NUBench training data, including configuration of the fine-tuning and instruction-based examples in Alpaca format (alpaca).
The YAML configuration provided in Listing 10 is specific to the LLaMA-3.1-8B model. Other models (e.g., Qwen or Mistral) can be fine-tuned similarly by modifying the model_name_or_path and template fields in the configuration file accordingly.
@@
Listing 10: Llama-3.1-8B_lora_sft.yaml
@@
3 "instruction": "Negate the sentence.",
4 "input": "Sentence: Eddie Vedder was born before Nam Woo-hyun.",
@@
8 "instruction": "Negate the sentence.",
9 "input": "Sentence: Halestorm is from Pennsylvania, while Say Anything is from California.",
@@
Listing 11: Sentence-Negation Pair dataset for training in alpaca format.
Appendix O Prompt Selection for In-Context Learning
We explore a range of prompt types for in-context learning, from minimal instructions to more detailed variants (zhao2021calibrate; li2023practical; wan2023better). Specifically, we evaluate three prompt styles:
  1. 1.
    Simple prompt: a minimal instruction, “Negate the sentence.”
  2. 2.
    Definition prompt: a concise description of the task, “Negate the main predicate of the main clause so that the proposition is logically reversed.”
  3. 3.
    Detail prompt: an extended instruction with explicit step-by-step guidelines on identifying the main predicate, preserving other sentence elements, handling antonyms, and applying negation consistently across logical operators (see Listing 12 for the full format).
"Reverse the truth value of the main predicate (verbal phrase) in the main clause, while preserving all other elements of the main clause unchanged.
1) Identify the main clause and its main verb (main predicate). Ignore subordinate clauses.
2) Preserve all other main-clause content.
3) Insert a negative particle such as "not" into the main verb, or replace it with a complementary antonym only if it forms an absolute binary (e.g., alive/dead, true/false, possible/impossible).
4) If the sentence contains multiple propositions connected by logical operators (e.g., and, or, conditional constructions), negate it in a way that reverses the entire proposition (e.g., A and B => not A or not B; If A then B => A and not B).
Sentence: {doc[’sentence’]}"
Listing 12: Detail prompt format.
Across both completion-based and option-selection evaluation settings, we found that the simple prompt consistently achieved the highest average performance in zero-shot and few-shot settings (See results in Appendix P.) While the definition-based and detail prompts occasionally provided more explicit guidance, they did not improve performance overall. Therefore, we adopt the simple prompt as our default setting in the main experiments. The complete prompt formats, including prompt–response structures for each evaluation setting, are provided in Listings 13 and 14, based on the simple prompt.
Prompt] "Negate the sentence.
Sentence: Chromosome 2 is the second-largest human chromosome, spanning more than 242 million base pairs and representing almost eight percent of the total DNA in human cells.
Negation:"
Response] "Chromosome 2 isnt the second-largest human chromosome, which measures more than 242 million base pairs and represents almost eight percent of the entire DNA in human cells."
Listing 13: Completion-based Format.
Prompt] "Given the following instruction and candidate answers, choose the single best answer.
Instruction: Negate the sentence.
Sentence: Chromosome 2 is the second-largest human chromosome, spanning more than 242 million base pairs and representing almost eight percent of the total DNA in human cells.
A. Chromosome 2 isnt the second-largest human chromosome, which measures more than 242 million base pairs and represents almost eight percent of the entire DNA in human cells.
B. Chromosome 2 is the second-largest human chromosome, which doesnt span more than 242 million base pairs or represent nearly eight percent of the whole DNA in human cells.
C. Chromosome 2 is the smallest human chromosome, spanning fewer than 50 million base pairs and representing less than two percent of the total DNA in human cells.\
D. Chromosome 2 is the second-biggest human chromosome, with over 242 million base pairs, making up nearly 8% of all DNA in human cells.
Your response should be one of A, B, C, D.
Only output the letter.
Answer:"
Response] "A"
Listing 14: Option-selection Format.
Appendix P Total Model Performance on Thunder-NUBench
zero-shot 1-shot 5-shot 10-shot After SFT
evaluation
setting
comple-
tion
option
comple-
tion
@@
comple-
tion
(±SD)
option
(±SD)
comple-
tion
option
gemma-2b 0.437 0.197
0.437
(±0.005)
0.248
(±0.005)
0.524
(±0.005)
0.276
(±0.007)
0.572
(±0.004)
0.267
(±0.011)
0.732 0.259
gemma-1.1-
2b-it
0.388 0.264
0.428
(±0.005)
0.347
(±0.004)
0.526
(±0.009)
0.357
(±0.008)
0.540
(±0.008)
0.336
(±0.004)
0.750 0.007
gemma-7b 0.454 0.388
0.465
(±0.001)
0.631
(±0.012)
0.561
(±0.002)
0.691
(±0.010)
0.620
(±0.010)
0.720
(±0.006)
0.797 0.722
gemma-1.1-
7b-it
0.663 0.645
0.689
(±0.009)
0.684
(±0.006)
0.730
(±0.007)
0.714
(±0.003)
0.759
(±0.007)
0.707
(±0.004)
0.786 0.733
Llama-3.2-3B 0.444 0.313
0.481
(±0.006)
0.417
(±0.011)
0.574
@@
(±0.009)
0.601
(±0.007)
0.501
(±0.009)
0.733 0.342
Llama-3.2-3B-
Instruct
0.472 0.530
0.487
(±0.009)
0.540
(±0.011)
0.557
(±0.006)
0.526
(±0.006)
0.587
(±0.004)
0.532
(±0.005)
0.745 0.604
Llama-3.1-8B 0.439 0.459
0.493
(±0.007)
0.549
(±0.012)
0.596
(±0.006)
0.635
(±0.008)
0.635
(±0.006)
0.667
(±0.011)
0.797 0.556
Llama-3.1-8B-
Instruct
0.512 0.744
0.571
(±0.007)
0.708
(±0.007)
0.646
(±0.007)
0.726
(±0.008)
0.675
(±0.004)
0.729
(±0.011)
0.756 0.747
Mistral-7B-v0.3 0.425 0.376
0.471
(±0.003)
0.603
(±0.008)
0.548
(±0.012)
0.702
(±0.010)
0.597
(±0.002)
0.720
(±0.013)
0.773 0.386
Mistral-7B-
Instruct-v0.3
0.619 0.642
0.635
(±0.007)
0.630
(±0.007)
0.664
(±0.005)
0.664
(±0.004)
0.690
(±0.008)
0.667
(±0.007)
0.765 0.650
Qwen2.5-3B 0.458 0.500
0.470
(±0.008)
0.538
(±0.008)
0.520
(±0.005)
0.560
(±0.008)
0.554
(±0.006)
0.567
(±0.009)
0.724 0.629
Qwen2.5-3B-
Instruct
0.534 0.620
0.599
(±0.007)
0.584
(±0.003)
0.633
(±0.009)
0.600
(±0.007)
0.656
(±0.007)
0.609
(±0.006)
0.700 0.726
Qwen2.5-7B 0.480 0.623
0.474
(±0.008)
0.608
(±0.007)
0.504
(±0.004)
0.622
(±0.007)
0.556
(±0.006)
0.629
(±0.003)
0.761 0.656
Qwen2.5-7B-
Instruct
0.542 0.676
0.558
(±0.007)
0.637
(±0.005)
0.620
(±0.004)
0.623
(±0.005)
0.659
(±0.005)
0.614
(±0.005)
0.718 0.793
gpt-4o-mini - 0.643 -
0.682
(±0.002)
0.687
(±0.005)
0.704
(±0.008)
gpt-4.1-mini - 0.797 -
0.822
(±0.005)
0.841
(±0.008)
0.846
(±0.005)
claude-3-5-
haiku-latest
- 0.722 -
0.505
(±0.016)
0.476
(±0.009)
0.633
(±0.004)
2-3B models
average
0.456 0.404 0.484 0.446 0.556 0.464 0.585 0.469 0.731 0.428
7-8B models
average
0.517 0.569 0.545 0.631 0.609 0.672 0.649 0.682 0.769 0.655
api models
average
- 0.721 - 0.670 - 0.668 - 0.728 - -
base models
average
0.448 0.408 0.470 0.513 0.547 0.564 0.591 0.582 0.760 0.507
instruct-tuned
models average
0.533 0.589 0.567 0.59 0.625 0.601 0.652 0.599 0.746 0.608
total average 0.491 0.538 0.518 0.573 0.586 0.598 0.622 0.615 0.753 0.558
Table 11: Zero-shot, few-shot, and SFT evaluation results on Thunder-NUBench with the simple prompt. SD denotes standard deviation across random seeds or runs. Few-shot results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and one, five, and ten demonstration examples (1-, 5-, 10-shot). Red text indicates the model with the highest performance in each column (excluding API models).
zero-shot 1-shot 5-shot 10-shot After SFT
evaluation
setting
comple-
tion
option
comple-
tion
(±SD)
option
(±SD)
comple-
tion
(±SD)
option
(±SD)
comple-
tion
(±SD)
option
(±SD)
comple-
tion
option
gemma-2b 0.443 0.176
0.427
(±0.002)
0.260
(±0.002)
0.515
(±0.005)
0.272
(±0.009)
0.560
(±0.007)
0.266
(±0.014)
0.720 0.259
gemma-1.1-
2b-it
0.402 0.263
0.436
(±0.004)
0.322
(±0.006)
0.542
(±0.006)
0.319
(±0.004)
0.570
(±0.005)
0.309
(±0.007)
0.753 0.334
gemma-7b 0.501 0.404
0.484
(±0.006)
0.568
(±0.009)
0.581
(±0.008)
0.638
(±0.008)
0.630
(±0.002)
0.675
(±0.006)
0.803 0.709
gemma-1.1-
7b-it
0.656 0.614
0.680
(±0.006)
0.627
(±0.009)
0.728
(±0.006)
0.657
(±0.006)
0.748
(±0.005)
0.668
(±0.006)
0.775 0.692
Llama-3.2-3B 0.416 0.324
0.471
(±0.006)
0.381
(±0.008)
0.561
(±0.007)
0.451
(±0.008)
0.590
(±0.008)
0.491
(±0.009)
0.754 0.282
Llama-3.2-3B-
Instruct
0.454 0.516
0.477
(±0.006)
0.464
(±0.010)
0.546
(±0.005)
0.452
(±0.008)
0.589
(±0.004)
0.450
(±0.006)
0.769 0.386
Llama-3.1-8B 0.462 0.477
0.499
(±0.005)
0.512
(±0.013)
0.598
(±0.006)
0.591
(±0.008)
0.645
(±0.005)
0.630
(±0.009)
0.817 0.608
Llama-3.1-8B-
Instruct
0.519 0.684
0.557
(±0.005)
0.640
(±0.004)
0.628
(±0.002)
0.663
(±0.004)
0.665
(±0.007)
0.681
(±0.014)
0.784 0.762
Mistral-7B-v0.3 0.446 0.297
0.475
(±0.006)
0.571
(±0.013)
0.560
(±0.011)
0.683
(±0.010)
0.608
(±0.002)
0.694
(±0.007)
0.781 0.427
Mistral-7B-
Instruct-v0.3
0.596 0.630
0.635
(±0.006)
0.622
(±0.007)
0.685
(±0.005)
0.630
(±0.011)
0.707
(±0.009)
0.632
(±0.007)
0.765 0.629
Qwen2.5-3B 0.427 0.531
0.474
(±0.007)
0.490
(±0.007)
0.540
(±0.003)
0.499
(±0.006)
0.574
(±0.007)
0.500
(±0.007)
0.722 0.648
Qwen2.5-3B-
Instruct
0.474 0.585
0.530
(±0.008)
0.511
(±0.006)
0.628
(±0.007)
0.512
(±0.006)
0.665
(±0.010)
0.513
(±0.006)
0.696 0.713
Qwen2.5-7B 0.454 0.628
0.483
(±0.009)
0.606
(±0.005)
0.541
(±0.007)
0.622
(±0.007)
0.587
(±0.002)
0.639
(±0.007)
0.745 0.706
Qwen2.5-7B-
Instruct
0.536 0.636
0.549
(±0.007)
0.634
(±0.009)
0.634
(±0.010)
0.634
(±0.008)
0.674
(±0.006)
0.636
(±0.004)
0.734 0.764
gpt-4o-mini - 0.621 -
0.659
(±0.008)
0.688
(±0.005)
0.700
(±0.009)
gpt-4.1-mini - 0.842 -
0.845
(±0.007)
0.857
(±0.005)
0.867
(±0.005)
claude-3-5-
haiku-latest
- 0.688 -
0.495
(±0.006)
0.470
(±0.008)
0.613
(±0.013)
2-3B models
average
0.436 0.399 0.469 0.405 0.555 0.418 0.591 0.422 0.736 0.437
7-8B models
average
0.521 0.546 0.545 0.598 0.619 0.640 0.658 0.657 0.775 0.662
api models
average
- 0.717 - 0.666 - 0.672 - 0.727 - -
base models
average
0.450 0.405 0.473 0.484 0.557 0.537 0.599 0.556 0.763 0.520
instruct-tuned
models average
0.520 0.561 0.552 0.546 0.627 0.552 0.660 0.556 0.754 0.611
total average 0.485 0.524 0.513 0.542 0.592 0.567 0.629 0.586 0.758 0.566
Table 12: Zero-shot, few-shot, and SFT evaluation results on Thunder-NUBench with the definition prompt. SD denotes standard deviation across random seeds or runs. Few-shot results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and one, five, and ten demonstration examples (1-, 5-, 10-shot). Red text indicates the model with the highest performance in each column (excluding API models).
zero-shot 1-shot 5-shot 10-shot After SFT
evaluation
setting
comple-
tion
option
comple-
tion
(±SD)
option
(±SD)
comple-
tion
(±SD)
option
(±SD)
comple-
tion
(±SD)
option
(±SD)
comple-
tion
option
gemma-2b 0.404 0.257
0.423
(±0.003)
0.253
(±0.005)
0.511
(±0.003)
0.257
(±0.012)
0.547
(±0.004)
0.145
(±0.009)
0.496 0.514
gemma-1.1-
2b-it
0.394 0.260
0.440
(±0.007)
0.282
(±0.005)
0.524
(±0.009)
0.241
(±0.007)
0.561
(±0.009)
0.242
(±0.009)
0.519 0.476
gemma-7b 0.462 0.410
0.472
(±0.005)
0.483
(±0.008)
0.564
(±0.004)
0.531
(±0.007)
0.607
(±0.010)
0.549
(±0.008)
0.731 0.635
gemma-1.1-
7b-it
0.588 0.472
0.653
(±0.008)
0.510
(±0.009)
0.708
(±0.009)
0.501
(±0.004)
0.728
(±0.006)
0.533
(±0.011)
0.725 0.616
Llama-3.2-3B 0.445 0.306
0.463
(±0.005)
0.317
(±0.011)
0.550
(±0.008)
0.365
(±0.010)
0.582
(±0.006)
0.407
(±0.008)
0.567 0.553
Llama-3.2-3B-
Instruct
0.454 0.429
0.440
(±0.006)
0.387
(±0.009)
0.501
(±0.007)
0.381
(±0.008)
0.560
(±0.005)
0.389
(±0.006)
0.586 0.595
Llama-3.1-8B 0.442 0.402
0.466
(±0.004)
0.394
(±0.008)
0.566
(±0.006)
0.462
(±0.012)
0.624
@@
0.544
(±0.003)
0.674 0.705
gpt-4o-mini - 0.595 -
0.620
(±0.014)
0.647
(±0.003)
0.669
(±0.006)
gpt-4.1-mini - 0.858 -
0.864
(±0.004)
0.876
(±0.003)
0.870
@@
claude-3-5-
haiku-latest
- 0.604 -
0.472